OpenStudy (anonymous):

Could someone please explain how to get answers a nd b below: it has to be in the form of ax+b than y''=0 y'=a y=ax+b -2a+4(ax+b)=2+3x 4ax-2a+4b=2+3x a=3/4 b=8/7

OpenStudy (anonymous):

differentiate it once and twice, u get 2 eqs, solve them and u get a and b values

OpenStudy (anonymous):

can you show me the working

OpenStudy (anonymous):

Is this given in the question : -2a+4(ax+b)=2+3x 4ax-2a+4b=2+3x or did u try to find these eqs??

OpenStudy (anonymous):

yeah its part of a differential equation question

OpenStudy (anonymous):

yeah thats correct, now differentiate the equations

OpenStudy (anonymous):

-2a+4(y)=2+3x ---> y=(2+3x+2a)/4 y' = 3/4 = a b is just replacing in ax+b = (2+3x+2a)/4